# LinkAlign: Scalable Schema Linking for Real-World Large-Scale Multi-Database Text-to-SQL

### v3

@@ Abstract
Yihan Wang1thanks:    Authors contributed equally.  Peiyu Liu2* 1Beihang University  2University of International Business and Economics yihan3123@gmail.comliupeiyustu@163.com
@@ 1 Introduction
Motivated by these factors, we propose LinkAlign, a novel framework that systematically addresses the challenges of schema linking in real-world environment. To address Challenge 1, our approach focuses first on (1) Retrieving Potential Database Schemas through multi-round semantic alignment by query rewriting. This step infers missing database schemas from retrieval results by leveraging the LLM’s reflective capabilities, then rewrites the query to align with the ground truth schema semantically. Then our approach centers on (2) isolating irrelevant schema information to reduce noise by response filtering. This step filters out the database noise from a set of candidates, minimizes interference from irrelevant schemas, and streamlines the downstream processing pipeline. To address Challenge 2, our approach directs efforts towards (3) extracting schemas for SQL generation through identifying tables and columns by schema parsing. This step scales schema linking to large-scale databases by introducing advanced reasoning-enhanced prompting techniques like multi-agent debate Chan et al. (2023) and chain-of-thought Wei et al. (2022). To balance efficiency and accuracy, we propose two complementary implementation paradigms: Pipeline and Agent. The pipeline mode executes each step via a process-fixed single LLM call, offering a streamlined, low-latency solution ideal for real-time database queries. In contrast, the agent mode performs multi-turn agent collaboration during inference, harnessing test-time computation to scale schema linking capabilities to databases with massive and complex schemas.
@@ 2 Error Analysis
Error 1: Target Database without Retrieval (Database Retrieval). This error indicates that the retrieved results do not include complete ground-truth database schemas, accounting for 23.6% of the failures. For example, the user intends to query "which semester the master and the bachelor both got enrolled in", but the retrieved results exclude $degree\_program$ table from the target database, which requires inference based on query semantics and commonsense knowledge. However, general vectorized retrieval approaches only return semantically similar results based on embedding distance, conflicting with the fact that user complex queries often misaligned with the ground-truth schema.
Error 2: Referring Irrelevant Databases (Database Retrieval). Unlike Error 1 that focuses on the missing target database, Error 2 centers on the irrelevant schema noise introduced by imprecise retrieval. This error indicates that the model refers to incorrect schemas from unrelated databases when generating SQL, accounting for 13.3% of the failures. For example, the user intends to query "The first name of students who have both cats and dogs". However,the generated SQL incorrectly infers $People.first\_name$ from an unrelated database while overlooking the ground truth $Student.f\_name$, even though both are successfully retrieved and fed into the model. Without isolating irrelevant database information, the model tends to mistakenly select seemingly more appropriate schemas from unrelated databases, leading to SQL execution failures.
@@
Error 3: Linking to the Wrong Tables (Schema Item Grounding). This error indicates that the generated SQL omits or misuses tables, accounting for 19.8% of the failures. For example, both the $student$ and $people$ tables have fields for $name$, but the model selects the latter incorrectly. In real-world scenarios, the models often overlook critical tables, which directly impairs the execution accuracy of the generated SQL Maamari et al. (2024).
Error 4: Linking to the Wrong Columns (Schema Item Grounding). This error indicates that the generated SQL omits or misuses fields despite referencing the ground truth table correctly, accounting for 11.6% of the failures. For example, the model may omit the join columns $pets.pet\_id$ and $has\_pet.pet\_id$ in the join operation of the correct SQL statement. Missing such critical columns directly incurs SQL execution failure.
@@ 3.3 Component Implementation Optimization
Align Semantics by Query Rewriting. Inspired by the reflective capabilities of LLMs demonstrated by Shinn et al., we introduce a semantic-enhanced retrieval approach based on retrieval feedback to achieve precise alignment with the ground-truth schema. Specifically, the Schema Auditor initiates by mapping the user queries into structured triplets (entities, attributes, and constraints). Next, it scrutinizes the retrieval results to infer missing schemas that may critical for accurate SQL generation (e.g., tables or fields for SELECT, JOIN, or WHERE clauses). This process culminates in an audit report that summarizes the parsed query, the inferred missing schemas, and the corresponding confidence levels. Subsequently, the Query Rewriter Agent leverages the comprehensive report to enhance the original query by disambiguating ambiguous expressions, supplementing semantic information for missing elements, and transforming the query into a template format optimized for text embedding models, ultimately improving retrieval recall for the ground-truth schema.
Reduce Noise by Response Filtering. When multiple candidate schemas exhibit minimal semantic differentiation, achieving consensus through multi-agent debate can significantly mitigate the risk of confusion. Inspired by this insight, we meticulously designed a multi-agent debate model comprising two distinct LLM agents: Data Analyst and Database Expert. Specifically, the Data Analyst evaluates the alignment between each database and the user query based on domain relevance and schema coverage completeness, then ranking them through corresponding comprehensive assessment. The highest ranked database is then selected from all candidates, with its schema and query context provided to the Database Expert. Subsequently, the Database Expert rigorously evaluates whether its provided database schema can satisfy the query requirements, validating the selection’s appropriateness through pseudo-SQL generation and determining whether to retain it. The debate follows a one-by-one strategy, i.e., starting with the data analyst, after which the two roles present their perspectives in turn. The debate ends when a predefined number of rounds is reached, and then a terminator outputs the consensus database as a final result.
@@ 4.1 Experiemental Setup
$\bullet$ Locate Accuracy (LA). Let $N$ denote the total number of test examples and $N_{a}$ the number of examples without Error 1 or Error 2. The LA is defined as $N_{a}$ / $N$, measuring the model’s ability to locate the target database accurately.
$\bullet$ Exact Matching (EM). Let $N_{e}$ denote the number of examples without Error 1 to 4. The EM score is defined as $N_{e}$ / $N$, measuring the model’s ability to perform precise schema linking.
@@ 5 Conclusion
6 Limitations
@@ Appendix A Related Work
Emerging Solutions and Challenges. Schema linking remains a challenge when handling large-scale schema and further complicated by the ambiguity in user queries. Approaches such as CHESS Talaei et al. (2024) and MCS-SQL Lee et al. (2024) try to address this by using multiple intricate prompts and sampling responses from LLMs on existing benchmarks such as Spider Yu et al. (2018) and Bird Li et al. (2024a). Although effective, they still require considerable computational resources and API costs, particularly in large-scale database scenarios.
@@ Appendix C Single-Prompt Pipeline Strategy
Appendix D Effectiveness in Single-Datbase Setting

### v4

@@ Abstract
Yihan Wang1,2thanks:   Authors contributed equally.  Peiyu Liu3*  Xin Yang1thanks:   Corresponding author. 1China Academy of Information and Communications Technology 2Renmin University of China 3University of International Business and Economics yihan3123@gmail.comliupeiyustu@163.comyangxincps@163.com
@@ 1 Introduction
Motivated by these factors, we propose LinkAlign, a novel framework that systematically addresses the challenges of schema linking in real-world environment. To address Challenge 1, our approach focuses first on (1) Retrieving Potential Database Schemas through multi-round semantic alignment by query rewriting. This step infers missing database schemas from retrieval results by leveraging the LLM’s reflective capabilities, then rewrites the query to align with the ground truth schema semantically. Then our approach centers on (2) isolating irrelevant schema information to reduce noise by response filtering. This step filters out the database noise from a set of candidates, minimizes interference from irrelevant schemas, and streamlines the downstream processing pipeline. To address Challenge 2, our approach directs efforts towards (3) extracting schemas for SQL generation through identifying tables and columns by schema parsing. This step scales schema linking to large-scale databases by introducing advanced reasoning-enhanced prompting techniques like multi-agent debate Chan et al. (2023); Pei et al. (2025) and chain-of-thought Wei et al. (2022). To balance efficiency and accuracy, we propose two complementary implementation paradigms: Pipeline and Agent. The pipeline mode executes each step via a process-fixed single LLM call, offering a streamlined, low-latency solution ideal for real-time database queries. In contrast, the agent mode performs multi-turn agent collaboration during inference, harnessing test-time computation to scale schema linking capabilities to databases with massive and complex schemas.
@@ 2 Error Analysis
Error 1: Target Database without Retrieval (Database Retrieval). This error indicates that the retrieved results do not include complete ground-truth database schemas, accounting for 23.6% of the failures. For example, the user intends to query "which semester the master and the bachelor both got enrolled in", but the retrieved results exclude degree_programdegree\_program table from the target database, which requires inference based on query semantics and commonsense knowledge. However, general vectorized retrieval approaches only return semantically similar results based on embedding distance, conflicting with the fact that user complex queries often misaligned with the ground-truth schema.
Error 2: Referring Irrelevant Databases (Database Retrieval). Unlike Error 1 that focuses on the missing target database, Error 2 centers on the irrelevant schema noise introduced by imprecise retrieval. This error indicates that the model refers to incorrect schemas from unrelated databases when generating SQL, accounting for 13.3% of the failures. For example, the user intends to query "The first name of students who have both cats and dogs". However,the generated SQL incorrectly infers People.first_namePeople.first\_name from an unrelated database while overlooking the ground truth Student.f_nameStudent.f\_name, even though both are successfully retrieved and fed into the model. Without isolating irrelevant database information, the model tends to mistakenly select seemingly more appropriate schemas from unrelated databases, leading to SQL execution failures.
@@
Error 3: Linking to the Wrong Tables (Schema Item Grounding). This error indicates that the generated SQL omits or misuses tables, accounting for 19.8% of the failures. For example, both the studentstudent and peoplepeople tables have fields for namename, but the model selects the latter incorrectly. In real-world scenarios, the models often overlook critical tables, which directly impairs the execution accuracy of the generated SQL Maamari et al. (2024).
Error 4: Linking to the Wrong Columns (Schema Item Grounding). This error indicates that the generated SQL omits or misuses fields despite referencing the ground truth table correctly, accounting for 11.6% of the failures. For example, the model may omit the join columns pets.pet_idpets.pet\_id and has_pet.pet_idhas\_pet.pet\_id in the join operation of the correct SQL statement. Missing such critical columns directly incurs SQL execution failure.
@@ 3.3 Component Implementation Optimization
Align Semantics by Query Rewriting. Inspired by the reflective capabilities of LLMs demonstrated by Shinn et al., we introduce a semantic-enhanced retrieval approach based on retrieval feedback to achieve precise alignment with the ground-truth schema. Specifically, the Schema Auditor initiates by mapping the user queries into structured triplets (entities, attributes, and constraints). Next, it scrutinizes the retrieval results to infer missing schemas that may critical for accurate SQL generation (e.g., tables or fields for SELECT, JOIN, or WHERE clauses). This process culminates in an audit report that summarizes the parsed query, the inferred missing schemas, and the corresponding confidence levels. Subsequently, the Query Rewriter Agent leverages the comprehensive report to enhance the original query by clarifying ambiguous expressions, supplementing semantic information for missing elements, and transforming the query into a template format optimized for text embedding models, ultimately improving retrieval recall for the ground-truth schema.
Reduce Noise by Response Filtering. When multiple candidate schemas exhibit minimal semantic differentiation, achieving consensus through multi-agent debate can significantly mitigate the risk of confusion. Inspired by this insight, we meticulously designed a multi-agent debate model comprising two distinct LLM agents: Data Analyst and Database Expert. Specifically, the Data Analyst evaluates the alignment between each database and the user query based on domain relevance and schema coverage completeness, then ranking them through corresponding comprehensive assessment. The highest ranked database is then selected from all candidates, with its schema and query context provided to the Database Expert. Subsequently, the Database Expert rigorously evaluates whether its provided database schema can satisfy the query requirements, validating the selection’s appropriateness and determining whether to retain it. The debate follows a one-by-one strategy, i.e., starting with the data analyst, after which the two roles present their perspectives in turn. The debate ends when a predefined number of rounds is reached, and then a terminator outputs the consensus database as a final result.
@@ 4.1 Experimental Setup
$\bullet$ Locate Accuracy (LA). Let $N$ denote the total number of test examples and $N_{a}$ the number of examples without Error 1 or Error 2. The LA is defined as $_{a} / NN$, measuring the model’s ability to locate the target database accurately.
$\bullet$ Exact Matching (EM). Let $N_{e}$ denote the number of examples without Error 1 to 4. The EM score is defined as $_{e} / NN$, measuring the model’s ability to perform precise schema linking.
@@ 5 Conclusion
Acknowledgments
This work was partially supported by National Natural Science Foundation of China under Grant No. 62506077.
@@ Appendix A Related Work
Emerging Solutions and Challenges. Schema linking remains a challenge when handling large-scale schema and further complicated by the ambiguity in user queries. Approaches such as CHESS Talaei et al. (2024) and MCS-SQL Lee et al. (2024) try to address this by using multiple intricate prompts and sampling responses from LLMs on existing benchmarks such as Spider Yu et al. (2018) and Bird Li et al. (2024a). To address challenges, some works focus on system robustness and mitigating model hallucinations. For instance, Solid-SQL Liu et al. (2024a) enhances robustness via a specialized pre-processing pipeline, while TA-SQL Qu et al. (2024) introduces a “Task Alignment” strategy to reduce hallucinations by reframing sub-tasks into more familiar problems for the model. Although effective, they still require considerable computational resources and API costs, particularly in large-scale database scenarios. An alternative line of research is to explore traditional model compression techniques—such as pruning Xia et al. (2024); Liu et al. (2025), quantization Lin et al. (2024); Liu et al. (2024b, c), or distillation Hsieh et al. (2023), to adapt LLMs into smaller Liu et al. (2021), task-specific variants. This may provide a cost-efficient and scalable direction for future work.
@@ Appendix C Single-Prompt Pipeline Strategy
Appendix D Effectiveness in Single-Database Setting
@@ Appendix F Supplementary Experimental Setup
Appendix G End-to-End Efficiency Analysis
To further assess LinkAlign’s efficiency, we conducted an additional evaluation on the Spider 2.0-Lite benchmark, which is recognized for its large-scale and realistic database settings. We randomly sampled 50 examples and adopted DeepSeek-V3 as the backbone to measure the end-to-end SQL generation efficiency of different baseline methods. As reported in Table 8, LinkAlign consistently improves the quality of SQL generation, without incurring notable overhead in runtime or token usage. In particular, the Pipeline mode exhibits clear efficiency gains over baselines that directly process the full schema (e.g., DIN-SQL), highlighting its effectiveness in balancing performance and cost.
Methods Avg. Time (s) Avg. Token
LinkAlign-Pipeline 127.6 8507.4
LinkAlign-Agent 183.5 12486.5
CHESS 457.8 21413.8
RSL-SQL 157.2 14713.4
DIN-SQL 146.3 8600.0
Table 8: End-to-end run time efficiency comparison. LinkAlign achieves improved schema linking without significant increases in runtime or token consumption.
Appendix H Limitations